OpenStudy (anonymous):

Question 8 There is no solution to the system below. x + 3y = 7 2x + 6y = 9 True False

OpenStudy (anonymous):

true

OpenStudy (anonymous):

[1 3 7 ] [2 6 9 ] inverse of [1 3] [2 6] is singular so no solution TRUE

OpenStudy (anonymous):

x=7-3y substitute gives 14-6y+6y=9 gives 14=9 which is false. So no solution. Answer True there is no solution. another method: multiply 1st line by 2 gives 2x+6y=14 so again 14=2x+6y=7 with same result.

OpenStudy (anonymous):

true

OpenStudy (anonymous):

[1 3 7 ] [2 6 9 ] cannot take inverse because matrix is singular so no solution TRUE